Methanol (CH3OH) is a volatile liquid used widely as a laboratory solvent.

Methanol has a standard boiling point of 64.8 °C and its enthalpy of vaporization is 37.6 kJ mol-1. What is its vapour pressure (in bar) at 38.1 °C?
1
Expert's answer
2019-12-03T08:24:26-0500

Use Clausius-Clapeyron equation


"ln\frac{P_1}{P_2} =\frac{\Delta H_{vap}}{R} (\frac{1}{T_2}-\frac{1}{T_1})"

"P_1 = 101325 Pa"

"T_1 (K) = 273 + 64.8 = 337.8 K"

"T_2(K) = 273 + 38.1 =311.1 K"


"ln\frac{101325}{P_2} = \frac{37600}{8.314} (\frac{1}{311.1}-\frac{1}{337.8})"

"P_2 = 32115 Pa"